# Gagan Bhatia ξ El Moatez Billah Nagoudi ξ Abdellah El Mekki ω Fakhraddin Alwajih ξ Muhammad Abdul-Mageedξ,ω,λ

### v2

@@ Abstract
We introduce Swan, a family of embedding models centred around the Arabic language, addressing both small-scale and large-scale use cases. Swan includes two variants: Swan-Small, based on ARBERTv2, and Swan-Large, built on ArMistral, a pretrained Arabic large language model. To evaluate these models, we propose ArabicMTEB, a comprehensive benchmark suite that assesses cross-lingual, multi-dialectal, multi-domain, and multi-cultural Arabic text embedding performance, covering eight diverse tasks and spanning 94 datasets. Swan-Large achieves state-of-the-art results, outperforming Multilingual-E5-large in most Arabic tasks, while the Swan-Small consistently surpasses Multilingual-E5-base. Our extensive evaluations demonstrate that Swan models are both dialectally and culturally aware, excelling across various Arabic domains while offering significant monetary efficiency. This work significantly advances the field of Arabic language modelling and provides valuable resources for future research and applications in Arabic natural language processing. Our models and benchmark will be made publicly accessible for research.
[Uncaptioned image] Swan and ArabicMTEB: Dialect-Aware, Arabic-Centric, Cross-Lingual, and Cross-Cultural Embedding Models and Benchmarks
Gagan Bhatia ξ El Moatez Billah Nagoudi ξ Abdellah El Mekki ω Fakhraddin Alwajih ξ Muhammad Abdul-Mageedξ,ω,λ ξ The University of British Columbia,      ω MBZUAI,      λ Invertible AI {gagan30@student.,muhammad.mageed@}ubc.ca
@@ 1 Introduction
Natural Language Processing (NLP) has seen rapid advancements in recent years, driven by groundbreaking developments in deep learning and the emergence of sophisticated distributed text representations such as word and sentence embeddings Devlin et al. (2018); Reimers and Gurevych (2019). These embeddings, which transform text into dense vectors, enable effective semantic understanding and are pivotal for enhancing performance across many downstream applications, including text classification, semantic search, and machine translation. Moreover, text embeddings have become fundamental to the success of large language models (LLMs) Touvron et al. (2023b, a); Jiang et al. (2023); Team et al. (2024), which are increasingly integrated into a variety of real-world systems and tools. One of the most promising applications of these embeddings is in the realm of Retrieval-Augmented Generation (RAG) Shao et al. (2023); rag (2023), where LLMs are augmented with information retrieval capabilities. In RAG-based systems, lightweight embedding models retrieve relevant information from large corpora, fed as context to models like ChatGPT OpenAI (2023) or GPT-4 OpenAI et al. (2024). This synergy between embeddings and LLMs has demonstrated significant improvements in both general-purpose tasks such as question answering Lin et al. (2023); rag (2023) and domain-specific applications Bhatia et al. (2024); Shi et al. (2023); Lin et al. (2023).
Figure 1: Details of ArabicMTEB
@@
Despite these advances, the predominant focus of current embedding models has been on English and Chinese, which limits their applicability to other languages. This is particularly true for Arabic, a language with rich morphology, diverse dialects, and unique syntactic structures, making it challenging to develop effective language representations Nagoudi et al. (2022); Huang et al. (2024). Existing multilingual models often fail to capture these complexities, leading to a suboptimal performance on Arabic NLP tasks Abdul-Mageed et al. (2020a); Elmadany et al. (2022). Addressing this limitation requires the development of Arabic-specific embedding models that are sensitive to the linguistic and cultural nuances of the language.
@@
In this work, we introduce Swan, a family of dialect-aware Arabic-centric cross-lingual cross-cultural embedding models designed to bridge this gap and push the boundaries of Arabic NLP. Our contributions are as follows:(1) We introduce Swan, a cutting-edge family of Arabic embedding models. This includes two variants: Swan-Small, based on ARBERTv2 Elmadany et al. (2022), and Swan-Large, built upon ArMistral, a further pretrained Arabic language model. (2) We present ArabicMTEB, a comprehensive evaluation benchmark for Arabic text. ArabicMTEB is designed to assess cross-lingual, multi-dialectal, multi-domain, and multi-cultural performance, spanning eight tasks and $94$ datasets. (3) Our larger model, Swan-Large, showcases state-of-the-art text embedding capabilities, surpassing Multilingual-E5-large Wang et al. (2024b) on most Arabic tasks. Additionally, our smaller, Swan-Small, consistently outperforms Multilingual-E5-base Wang et al. (2024b) on most Arabic tasks. (4) Through rigorous benchmarking, we demonstrate that Swan models are not only dialectally and culturally aware but also excel across diverse Arabic domains while maintaining a significantly lower monetary cost.
@@
The rest of the paper is organized as follows: In Section 2, we review related work with a particular emphasis on Arabic text embedding models, their applications and challenges. We present our approach to model training Swan models in Section 3. Section 4 outlines how we built our benchmark dataset, ArabicMTEB. Section 5 is about our experiments and model analysis. Finally, we conclude in Section 6.
@@ 2 Related Work
Benchmark Language Tasks Datasets #Tasks CRTR Arabic Culture/Domains
@@
Polish-MTEB Poświata et al. (2024) Polish RTR, STS, PairCLF, CLF, CLR 26 5 $\times$ $\times$
@@
Scand. MTEB Enevoldsen et al. (2024) Danish RTR, CLF, BTM, CLR 26 4 $\times$ $\times$
Norwegian $\times$ $\times$
@@
Table 1: Comparison of Various Text Embedding benchmarks proposed in the literature across the different covered task clusters. RTR: Retrieval, STS: Semantic Textual Similarity, PairCLF: Pair Classification, CLF: Classification, CLR: Clustering, RRK: Reranking, BTM: BitextMining, CRTR: Crosslingual Retrieval.
@@
Multilingual Text Embeddings Models. With the need for language-agnostic embeddings growing, multilingual models like LASER Artetxe and Schwenk (2019) and LaBSE Feng et al. (2022) were developed using BiLSTM and Transformer encoders, respectively. Building on this, the Multilingual-E5 Wang et al. (2024c) series extends the E5 architecture to support diverse languages using multilingual text pairs and synthetic data. GRIT Muennighoff et al. (2024) further unifies generative and embedding tasks within a single model. Newer models like ColBERT-XM Louis et al. (2024) and Gecko Lee et al. (2024) refine multilingual embeddings through modular and distilled architectures.
Arabic-Specific Models. Despite progress in Arabic NLP, models are not optimized for Arabic text embedding and retrieval. Efforts like ARBERT Abdul-Mageed et al. (2021a), ORCA Elmadany et al. (2023), and AraMus Alghamdi et al. (2023) have focused on encoding and generation but are not tailored for sentence-level embeddings. While language-agnostic models such as LASER and Multilingual-E5 include Arabic in their training data, they may not fully capture its linguistic intricacies. To address this, Nacar and Koubaa (2024) introduced models and training datasets to improve semantic similarity performance for Arabic.
@@ 3.1 Training Data
Monolingual Arabic Human ORCA-MSA Sentence 378K
@@
Synthetic Synth-MSA Paragraph 100K
@@
Crosslingual Arabic to 15 Langs Human MMARCO Sentence 3M
Arabic to 6 Langs XOR-TyDi 20.5K
Multilingual 11 Langs Human Mr-Tydi Sentence 49K
16 Langs Miracl 343K
@@
(a) Positive and hard negative generation
Figure 2: Methodology to generate our synthetic data.
MSA Datasets. We focus on two sources: (i) Human-generated data: Composed from ORCA Elmadany et al. (2023) and mMARCO Bonifacio et al. (2021). ORCA is a compilation of labelled datasets with tasks such as semantic text similarity (STS), sentence classification, text classification, natural language inference (NLI), and question answering. We use all the training sets from ORCA, encompassing $60$ different datasets. mMARCO-ar is the translated version of MARCO, which is a human-generated dataset Bajaj et al. (2018). Both of these datasets are cleaned up and de-duplicated using Polydedupe  Bhatia (2023), which is further described in Appendix D. (ii) Synthetically-generated data: To augment our MSA training data for retrieval tasks, we use Command R+ Cohere For AI (2024) to generate high-quality synthetic data. [^1] We performed various in-house evaluations comparing multiple models. Command R+ was chosen as it is open-source and efficient in generating Arabic varieties (standard and dialectal). The generation methodology was inspired by Wang et al. (2024a). We first generate tasks closely related to real-world usage using the model, and then we use the procedure shown in Figure 2 to generate our synthetic dataset. We generate $100k$ in general MSA data and $5k$ in instances for specific domains like finance, news, medicine, and legal for a total of $120k$ MSA instances.
Dialectal Arabic Datasets. Similar to the MSA datasets, we focus on two sources: (i) Human-generated data: We use publicly available dialectal Arabic data, which primarily covers Gulf, Egyptian, Moroccan, and Levantine varieties of Arabic Elmadany et al. (2022); Nagoudi et al. (2023b); Alwajih et al. (2024); Abdul-Mageed et al. (2020b, 2021c, 2022, 2023, 2018, c); Keleg et al. (2023); Keleg and Magdy (2023); Zaidan and Callison-Burch (2014); Bouamor et al. (2018). The total number of samples is $122K$. (ii) Synthetically-generated data: As most human-generated dialectal data comes from noisy environments such as social media, it often results in short texts of low quality. Thus, we use Command-R+ to generate paragraph-based synthetic data for Egyptian and Moroccan dialects to improve the performance of our models on dialectal Arabic. We generated $15k$ dialectal instances using the same methodology as our synthetic MSA datasets described above.
Cross-Lingual & Multilingual Datasets. To make our model adapted for cross-lingual and multilingual scenarios, we include the mMARCO dataset, which comprises translations of the MS MARCO dataset into $15$ languages Bonifacio et al. (2021). To ensure that documents correspond accurately to their queries in different languages, we utilize specific IDs. We create $100k$ samples for each cross-lingual pair and shuffle the IDs to prevent repetition, thus guaranteeing that unique data samples are employed for each language. We utilize the MIRACL Zhang et al. (2022), XOR-TyDI Asai et al. (2021), and Mr.TyDi Zhang et al. (2021) datasets as our crosslingual and multilingual resources.
@@ 3.2 Training Strategy
For Swan, we consider two models: Swan-Small and Swan-Large. The choice of training two models with different sizes, Swan-Small and Swan-Large, is driven by the need to balance performance and computational efficiency. Swan-Small is designed to cater to scenarios where lower computational resources are available or when a lightweight model is preferred for deployment on edge devices. In contrast, Swan-Large is intended for settings where achieving state-of-the-art performance is paramount, leveraging a larger parameter size to better capture the complexities of the Arabic language and its semantics.
@@
Data Preprocessing. We incorporated human-generated and synthetic datasets into our training pipeline to ensure robust performance across various dialects and cultural contexts. Initially, the model was trained on MSA datasets and then fine-tuned on dialectal datasets. This two-step approach ensures that both MSA and dialectal varieties are well represented, promoting better generalization across the full spectrum of Arabic language varieties. Our dataset was constructed with a query format, including positive and negative samples.
@@
Swan-Small. Built upon ARBERTv2 Abdul-Mageed et al. (2021b), a powerful BERT-based model for the Arabic language. Here, our model is trained using the InfoNCE loss van den Oord et al. (2019a), which maximizes the similarity between related sentences while minimizing the similarity between unrelated sentences. The model was trained for five epochs on the entire dataset with a learning rate of $5e^{-6}$ and a batch size of 128, incorporating fifteen hard negatives. Swan-Small has $164M$ parameters and a dimension size of 768.
@@
Swan-Large. Swan-Large is based on ArMistral-7B, an in-house further pretrained version of Mistral-7B Jiang et al. (2023) (Further details about ArMistral can be found in Appendix A). To train Swan-Large, we use LoRA Hu et al. (2021) for parameter efficient training and InfoNCE loss van den Oord et al. (2019a); Ma et al. (2023) for optimization. The model was trained for three epochs on the entire dataset with a learning rate of $5e^{-6}$ and a batch size of 128, incorporating seven hard negatives. Swan-Large has $7.2B$ parameters and a dimension size of 4096. More details about data preprocessing and training process can be found in Appendix B.
@@
Inclusion of Hard-Negatives To enhance the model’s performance, it is crucial to use negative documents closely aligned with the query’s context Karpukhin et al. (2020). This method allows us to observe the impact of introducing more challenging or "hard" negatives into the training process. We only generate hard negatives for the Arabic subset of our training data from Section 3.1. We found that using $15$ hard negatives for Swan-small yields the best performance, whereas for bigger model Swan-large, the model overfits a more significant number of hard negatives, and $7$ gives us the best performance. (Please see Appendix H for more details about the hard negatives exploration.)

### v3

# Dialect-Aware, Arabic-Centric, Cross-Lingual, and Cross-Cultural Embedding Models and Benchmarks

@@ Abstract
We introduce Swan, a family of embedding models centered on Arabic, designed for both small-scale and large-scale applications. Swan comprises two variants: Swan-Small, built on ARBERTv2, and Swan-Large, based on ArMistral, a pretrained Arabic large language model. To evaluate our models, we propose a comprehensive benchmark suite, dubbed ArabicMTEB, that assesses cross-lingual, multi-dialectal, multi-domain, and multi-cultural Arabic text embedding performance. ArabicMTEBcovers eight diverse tasks sourced from 94 datasets. Swan-Large achieves state-of-the-art results, outperforming Multilingual-E5-large in most Arabic tasks, while Swan-Small consistently surpasses Multilingual-E5-base. Our extensive evaluations show that Swan models are both dialectally and culturally aware, achieving strong performance across diverse Arabic domains while maintaining significant cost efficiency. This work significantly advances the field of Arabic language modelling and provides valuable resources for future research and applications in Arabic NLP. Our models and benchmark are available at our GitHub page: https://github.com/UBC-NLP/swan.
[Uncaptioned image] Swan and ArabicMTEB:
Dialect-Aware, Arabic-Centric, Cross-Lingual, and Cross-Cultural Embedding Models and Benchmarks
Gagan Bhatia ξ El Moatez Billah Nagoudi ξ,γ Abdellah El Mekki ω Fakhraddin Alwajih ξ Muhammad Abdul-Mageedξ,ω,λ ξ The University of British Columbia,      ω MBZUAI,       γ PSU,       λ Invertible AI {gagan30@student.,muhammad.mageed@}ubc.ca
@@ 1 Introduction
NLP has seen rapid advancements in recent years, driven by groundbreaking developments in deep learning and the emergence of sophisticated distributed text representations such as word and sentence embeddings Devlin et al. (2018); Reimers and Gurevych (2019). These embeddings, which transform text into dense vectors, enable effective semantic understanding and are pivotal for enhancing performance across many downstream applications, including text classification, semantic search, and machine translation. Moreover, text embeddings have become fundamental to the success of large language models (LLMs) Touvron et al. (2023); Jiang et al. (2023); Gemma-Team et al. (2024), which are increasingly integrated into a variety of real-world systems and tools. One of the most promising applications of these embeddings is in the realm of Retrieval-Augmented Generation (RAG) Shao et al. (2023); rag (2023), where LLMs are augmented with information retrieval capabilities. In RAG-based systems, lightweight embedding models retrieve relevant information from large corpora, fed as context to models like ChatGPT OpenAI (2023) or GPT-4 OpenAI et al. (2024). This synergy between embeddings and LLMs has demonstrated significant improvements in both general-purpose tasks such as question answering Lin et al. (2023); rag (2023) and domain-specific applications Bhatia et al. (2024); Shi et al. (2023); Lin et al. (2023).
Figure 1: Overview of our ArabicMTEB benchmark tasks, covering clustering, retrieval, reranking, classification, semantic similarity, pair classification, cross-lingual retrieval, and bitext mining.
@@
Despite these advances, the predominant focus of current embedding models has been on English and Chinese, which limits their applicability to other languages. This is particularly true for Arabic, a collection of languages, language varieties, and diverse dialects with rich morphology Abdul-Mageed et al. (2023a, 2024a), making it challenging to develop effective language representations Nagoudi et al. (2022); Huang et al. (2024). Existing multilingual models often fail to capture these complexities, leading to a suboptimal performance on Arabic NLP tasks Abdul-Mageed et al. (2020a); Elmadany et al. (2022). Addressing this limitation requires the development of Arabic-specific embedding models that are sensitive to the linguistic and cultural nuances of Arabic.
@@
In this work, we introduce Swan, a family of dialect-aware, Arabic-centric, cross-lingual, and cross-cultural embedding models designed to bridge this gap and push the boundaries of Arabic NLP. Our contributions are as follows:(1) We introduce Swan, a cutting-edge family of Arabic embedding models. This includes two variants: Swan-Small, based on ARBERTv2 Elmadany et al. (2022), and Swan-Large, built upon ArMistral, a further pretrained Arabic language model. (2) We present ArabicMTEB, a comprehensive evaluation benchmark for Arabic text. ArabicMTEB is designed to assess cross-lingual, multi-dialectal, multi-domain, and multi-cultural performance, spanning eight tasks and $94$ datasets. Figure 1 provides an overview of ArabicMTEB. (3) Our larger model, Swan-Large, showcases state-of-the-art text embedding capabilities, surpassing Multilingual-E5-large Wang et al. (2024b) on most Arabic tasks. Similarly, our smaller, Swan-Small, consistently outperforms Multilingual-E5-base Wang et al. (2024b) on most Arabic tasks. (4) Through rigorous benchmarking, we demonstrate that Swan models are not only dialectally and culturally aware, but also excel across diverse Arabic domains while maintaining a significantly lower monetary cost.
@@
The rest of the paper is organized as follows: In Section 2, we review related work with a particular emphasis on Arabic text embedding models, their applications and challenges. We present our approach to model training of Swan models in Section 3. Section 4 outlines how we built our benchmark dataset, ArabicMTEB. Section 5 is about our experiments and model analysis. We conclude in Section 6.
@@ 2 Related Work
Benchmark Lang Tasks Datasets Tasks CRTR Ar Cul/Dom
@@
Polish Poświata et al. (2024) Polish RTR, STS, PairCLF, CLF, CLR 26 5 $\times$ $\times$
@@
Scand. Enevoldsen et al. (2024) Danish RTR, CLF, BTM, CLR 26 4 $\times$ $\times$
Norweg. $\times$ $\times$
@@
Table 1: Comparison of various text embedding benchmarks proposed in the literature across the different covered task clusters. RTR: retrieval, STS: semantic textual similarity, PairCLF: pair classification, CLF: classification, CLR: clustering, RRK: reranking, BTM: bitext mining, CRTR: cross-lingual retrieval.
@@
Multilingual Text Embedding Models. With the need for language-agnostic embeddings growing, multilingual models such as LASER Artetxe and Schwenk (2019) and LaBSE Feng et al. (2022) were developed using BiLSTM and Transformer encoders, respectively. Building on this, the Multilingual-E5 Wang et al. (2024c) series extends the E5 architecture to support diverse languages using multilingual text pairs and synthetic data. GRIT Muennighoff et al. (2024) further unifies generative and embedding tasks within a single model. Newer models such as ColBERT-XM Louis et al. (2024) and Gecko Lee et al. (2024) refine multilingual embeddings through modular and distilled architectures.
Arabic-Specific Models. Despite progress in Arabic NLP, existing models are not optimized for Arabic text embedding and retrieval. Efforts like ARBERT Abdul-Mageed et al. (2021a) and AraMus Alghamdi et al. (2023) have focused on encoding and generation but are not tailored for sentence-level embeddings. While language-agnostic models such as LASER and Multilingual-E5 include Arabic in their training data, they may not fully capture its linguistic intricacies and diversity. To address this, Nacar and Koubaa (2024) introduced models and training datasets to improve semantic similarity performance for Arabic.
@@ 3.1 Training Data
Figure 2: Methodology to generate our synthetic data.
MSA Datasets. We focus on two sources: (i) Human-generated data: Composed from ORCA Elmadany et al. (2023) and mMARCO Bonifacio et al. (2021). ORCA is a compilation of labelled datasets with tasks such as semantic text similarity (STS), sentence classification, text classification, natural language inference (NLI), and question answering. We use all the training sets from ORCA, encompassing $60$ different datasets. mMARCO-ar is the translated version of MARCO, which is a human-generated dataset Bajaj et al. (2018). Both of these datasets are cleaned up and de-duplicated using Polydedupe  Bhatia (2023), [^1] https://github.com/gagan3012/PolyDeDupe which is further described in Appendix C. (ii) Synthetically-generated data: To augment our MSA training data for retrieval tasks, we use Command R+ Cohere For AI (2024) to generate high-quality synthetic data. [^2] We performed various in-house evaluations comparing multiple models. Command R+ was chosen as it is open-source and efficient in generating Arabic varieties (standard and dialectal). The generation methodology is inspired by Wang et al. (2024a), and we employ the procedure shown in Figure 2 to generate our synthetic dataset. We generate $100k$ in general MSA data and $5k$ in instances for specific domains such as finance, news, medicine, and legal for a total of $120k$ MSA instances.
@@
Monoling Ar Human ORCA-MSA Sent 378K
@@
Synthetic Synth-MSA Parag 100K
@@
Crossling Ar to 15 lg Human MMARCO Sent 3M
Ar to 6 lg XOR-TyDi 20.5K
Multiling 11 lg Human Mr-Tydi Sent 49K
16 lg Miracl 343K
@@
Dialectal Arabic Datasets. Similar to the MSA datasets, we focus on two sources: (i) Human-generated data: We use publicly available dialectal Arabic data, which primarily covers Gulf, Egyptian, Moroccan, and Levantine varieties of Arabic Elmadany et al. (2022); Nagoudi et al. (2023b); Alwajih et al. (2024); Abdul-Mageed et al. (2020b, 2021c, 2022, 2023b, 2018, c); Keleg et al. (2023); Keleg and Magdy (2023); Zaidan and Callison-Burch (2014); Bouamor et al. (2018). The total number of samples is $122K$. (ii) Synthetically-generated data: As most human-generated dialectal data comes from noisy environments such as social media, it often results in short texts of low quality. Thus, we use Command-R+ to generate paragraph-based synthetic data for Egyptian and Moroccan dialects to improve the performance of our models on dialectal Arabic. We generated $15k$ dialectal instances using the same methodology as our synthetic MSA datasets described above.
Cross-Lingual & Multilingual Datasets. To adapt our model for cross-lingual and multilingual scenarios, we incorporate the mMARCO dataset, which provides translations of the MS MARCO dataset into 15 languages Bonifacio et al. (2021). To ensure that documents correspond accurately to their queries in different languages, we utilize specific IDs. We create $100k$ samples for each cross-lingual pair and shuffle the IDs to prevent repetition, thus guaranteeing that unique data samples are used for each language. We utilize the MIRACL Zhang et al. (2022), XOR-TyDI Asai et al. (2021), and Mr.TyDi Zhang et al. (2021) datasets as our crosslingual and multilingual resources.
@@ 3.2 Training Strategy
For Swan, we consider two models: Swan-Small  and Swan-Large. The choice of training two models with different sizes is driven by the need to balance performance and computational efficiency. Swan-Small is designed to cater to scenarios where lower computational resources are available or when a lightweight model is preferred for deployment on edge devices. In contrast, Swan-Large is intended for settings where achieving SoTA performance is paramount, leveraging a larger parameter size to better capture the nuances of Arabic.
@@
Data Preprocessing. We incorporate human-generated and synthetic datasets into our training pipeline to ensure robust performance across various dialects and cultural contexts. We first train on MSA datasets, followedby fine-tuning on dialectal datasets. This two-step approach ensures that both MSA and dialectal varieties are well represented, promoting better generalization across the full spectrum of Arabic varieties. Our dataset is constructed with a query format, including positive and negative samples.
@@
Swan-Small. Built upon ARBERTv2 Abdul-Mageed et al. (2021b), a powerful BERT-based model for the Arabic language. Here, our model is trained using the InfoNCE loss van den Oord et al. (2019), which maximizes the similarity between related sentences while minimizing the similarity between unrelated sentences. The model is trained for five epochs on the entire dataset with a learning rate of $5e^{-6}$ and a batch size of 128, incorporating 15 hard negatives. Swan-Small has $164$M parameters and a dimension size of $768$.
@@
Swan-Large. Swan-Large is based on ArMistral-7B, an in-house further pretrained version of Mistral-7B Jiang et al. (2023) [^3] Further details about ArMistral can be found in Appendix A.. To train Swan-Large, we use LoRA Hu et al. (2021) for parameter efficient training and InfoNCE loss for optimization. We train the model for three epochs on the entire dataset with a learning rate of $5e^{-6}$ and a batch size of $128$, incorporating seven hard negatives. Swan-Large has $7.2B$ parameters and a dimension size of $4,096$.
3.3 Training Methodology
Here, “{task_instruction}” refers to a one-sentence description of the embedding task taken from Table 12, which outlines the instructions for different tasks. Using a pretrained LLM, we append an [EOS] token at the end of both the modified query and the document. These are then fed into the LLM to extract embeddings $\mathbf{h}_{q^{+}_{\text{inst}}}$ and $\mathbf{h}_{d^{+}}$ from the vector at the last [EOS] layer. Again, training of the embedding model is conducted using the InfoNCE loss function, which is widely recognized for its effectiveness in learning high-quality embeddings. The objective is minimized using the following formulation:
3.4 Inclusion of Hard-Negatives
@@
To enhance the model’s performance, it is crucial to use negative documents closely aligned with the query’s context Karpukhin et al. (2020); Khondaker et al. (2022). This method allows us to observe the impact of introducing more challenging or "hard" negatives into the training process. We only generate hard negatives for the Arabic subset of our training data from Section 3.1. We found that using $15$ hard negatives for Swan-Small yields the best performance, whereas for our bigger model, Swan-Large, the model overfits a more significant number of hard negatives, and $7$ gives us the best performance.
Impact of Hard Negatives. Hard negatives in contrastive learning are examples that closely resemble correct instances but are ultimately incorrect. Their inclusion encourages the model to learn finer-grained distinctions, improving its ability to differentiate between similar but distinct classes. The process involves converting all documents into a vector form within the embedding space. Subsequently, these document embeddings are compared using the cosine similarity score to establish their relevance to the query. Once all documents are scored, they are sorted by their similarity to the query. The top-ranked document is typically the positive example, while the rest are potential negatives. Our experiments assess the impact of varying the hard negatives used while training our models, Swan-Large and Swan-Small. We train each model with different quantities of hard negatives. Namely, we experiment with using hard negatives values from the set {1, 3, 7, 15, 31} per training instance. Swan-Small achieves its highest performance of 56.25 with 15 hard negatives. The model exhibits a general upward trend as the number of hard negatives increases, peaking at 15 before slightly declining at 31. This pattern suggests that while additional hard negatives initially enhance learning by introducing valuable challenges, excessive complexity may lead to diminishing returns, ultimately hindering further improvement. Swan-Large achieves its peak performance of 60.42 when trained with seven hard negatives, suggesting an optimal balance that enhances learning without overloading the model. Notably, increasing the number of hard negatives beyond this point does not lead to further gains, indicating a threshold where additional complexity ceases to improve learning outcomes.
Model (HN) 1 3 7 15 31
Swan-Small 48.84 52.19 54.13 56.25 51.93
Swan-Large 59.48 59.35 60.42 59.44 59.83
Table 3: Impact of number of Hard Negatives (HN).